HUBBLE SPACE TELESCOPE DAILY REPORT #1999 PERIOD COVERED: 0000Z (UTC) 11/05/97 - 0000Z (UTC) 11/06/97 Daily Status Report as of 310/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7621 (WF/PC-2 Cycle 7 Supplemental Darks PT1/3) The WF/PC-2 was used to make supplemental dark calibration images. This set of calibrations is intended to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed STIS/CCD 7802 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations to obtain daily darks to update the acquisition bad pixel table. These darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Nine Sets of STIS/CCD 7782 (STIS Non-Scripted Parallel Archive Proposal - Half Orbit) The Space Telescope Imaging Spectrograph (CCD) was used for a non-scripted parallel archive proposal (one-half orbit). The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Six Sets of WF/PC-2 7786 (Pure Parallel Archive Survey - Continuation) The WF/PC-2 was used to make observations as part of an archival survey program in which the images were taken in parallel with the prime science observations. These WF/PC-2 parallel observations were obtained in the ultraviolet in a search for interesting objects. The observations were completed as planned, and no problems were reported. 1.5 Completed Three Sets of STIS/CCD 7783 (STIS Parallel Science) The Space Telescope Imaging Spectrograph (CCD) was used to make observations taken in parallel with the prime science. The observations were completed, and no problems were reported. 1.6 Completed NIC/1/2/3 7447 (The Cosmological Evolution of Quasar Host Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of galaxy 4C02.54 as part of a survey to study the cosmological evolution of galaxies which host quasars. Other point spread function observations were taken to test the calibration of the cameras. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed STIS/CCD/MA1 7432 (Morphological and Spectral Evolution of Galaxies) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make observations to study the morphological and spectral evolution of the galaxy 3C17. A library of ultraviolet images of normal galaxies at low redshifts is essential to our understanding of the properties and evolution of the high redshift universe. The observations were completed as planned, and no problems were reported. 1.8 Completed NIC/1/2/3 7726 (NICMOS One Orbit Parallel) The Near Infrared Camera (Cameras One, Two and Three) was used to make a one orbit parallel calibration observations. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed WF/PC-2 7625 (WF/PC-2 Cycle 7 Earth Flats) The WF/PC-2 was used to make a series of Earth flat calibration images. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The observations were completed as planned, and no anomalies were reported. 1.10 Completed NIC/1/2/3 7729 (NICMOS Short One Orbit Parallel) The Near Infrared Camera (Cameras One, Two, and Three) was used to obtain "JHK" imaging of sky regions while another instrument was prime. These observations were part of the short NICMOS one-orbit parallel program. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The observations were completed as scheduled, and no problems were reported. 1.12 Completed NIC/2 7331 (Near-IR Properties of the Bulges of Spiral Galaxies) The Near Infrared Camera (Camera Two) was used to make infrared observations of the bulge in spiral galaxy NGC-7259. Near-IR data at HST resolution are essential for correcting for the effect of dust and will enable us to derive robust nuclear properties for the sample. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed NIC/1/2/3 7608 (NICMOS Focus Monitoring) The Near Infrared Cameras (Cameras One, Two, and Three) were used to determine and monitor the optimal focus and tilt settings for all three NICMOS cameras by making image of cluster NGC-3603. This proposal is built to run NIC1/2/3 sweeps on a 2-week basis. The observations completed without incident with no reported anomalies. 1.14 Completed WF/PC-2 7431 (Snapshot Survey of Microlensed Source Stars) The WF/PC-2 was used to make snapshot survey images of microlensed source star Macho-95-2. Gravitational microlensing surveys have revealed an excess of lensing events over theoretical predictions toward both the Large Magellanic Cloud and the Galactic bulge. The observations were completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 As described in HSTAR 6234, the acquisition at 310/0533Z defaulted to fine lock on FGS-3 only. This may have affected WF/PC 7430, which will appear in tomorrow's report. Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled : 15 Successful : 15 2.3 Operations Notes: Per ROP NS-3, the NSSC-1 status buffer was dumped and reset at 310/0631Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH